# Evaluating many-body stabilizer Rényi entropy by sampling reduced Pauli strings: singularities, volume law, and nonlocal magic

### v2

@@ Abstract
We present a novel quantum Monte Carlo method for evaluating the $\alpha$-stabilizer Rényi entropy (SRE) for any integer $\alpha\geq 2$. By interpreting $\alpha$-SRE as partition function ratios, we eliminate the sign problem in the imaginary-time path integral by sampling reduced Pauli strings within a reduced configuration space, which enables efficient classical computations of $\alpha$-SRE and its derivatives to explore magic in previously inaccessible 2D/higher-dimensional systems. Physically, we first separate the free energy contribution in $2$-SRE. At quantum critical points in 1D/2D transverse field Ising (TFI) models, we reveal nontrivial singularities associated with the characteristic function contribution, directly tied to magic. Their interplay leads to complicated behaviors of $2$-SRE, avoiding extrema at critical points generally. In contrast, by analyzing the volume-law correction of magic, which represents nonlocal magic residing in correlations, we find that its discontinuity is bound to critical properties and would be more useful than the full-state magic. Finally, we verify that $2$-SRE fails to characterize magic in mixed states (e.g. Gibbs states), yielding nonphysical results. This work provides a powerful tool for exploring the roles of magic in large-scale many-body systems, and reveals intrinsic relation between magic and many-body physics.
@@ I Introduction
Charactering many-body behaviors in quantum systems has been a central task since the last century [1, 2]. Despite the locality of interactions in most quantum matters, the intricate correlations that give rise to rich and exotic collective phenomena present formidable theoretical and computational hurdles for physicists. During the past decades, remarkable headways are made by introducing the theory of quantum information into many-body physics. Quantum entanglement, with its quantum nature and as probably the most representative quantum resource [3], has turned out to relate closely to phase transitions, conformal field theories and topological orders [4, 5, 6].
In recent years, another quantum resource called magic or non-stabilzierness, which is closely associated with the stabilizer formalism [7, 8], has also entered the field of vision in many-body physics [9, 10, 11]. According to the celebrated Gottesman-Knill theorem [12, 13, 14], magic, rather than merely entanglement, is what necessarily enables quantum computation to outperform classical computation, as simulating a highly entangled stabilizer state requires only polynomial resources on a classical computer with Clifford protocols.
@@
The contributions of this work span both algorithmic and physical levels. At the algorithmic level, we introduce the first unbiased Quantum Monte Carlo (QMC) algorithm capable of computing both the $\alpha$-SRE and its derivatives in large-scalce and high-dimensional quantum many-body systems within polynomial time, as long as the Hamiltonian is sign-problem free. In fact, over the past years, several classical numerical tools have been developed to calculate the $\alpha$-SRE including the methods with matrix product states (MPS) and tree tensor networks [16, 17, 18, 11, 19, 35]. However, these methods face limitations in two or higher dimensions and finite-temperature cases. A recent hybrid algorithm proposed in Ref. [36] addresses this issue by expressing the $\alpha$-SRE using the language of tensor network (TN) and then applying non-equilibrium QMC sampling [37]. However, their algorithm computes $\alpha$-SRE at only a single point per run and is incapable of extracting derivative information, yielding results with limited physical insight.
Technically, we prove that the QMC simulations of the $\alpha$-SRE, which relates to some generalized partition function, can be restricted in the reduced configuration space by sampling reduced Pauli strings (both terms are defined in Sec. III). Using this algorithm, it is straightforward to estimate the values and derivatives of $\alpha$-SRE by considering the partition function difference, in conjunction with Monte Carlo techniques including but not restricted to: the reweight-annealing method (importance sampling) [38, 39, 40, 41, 42, 43, 44, 45, 46], thermodynamic integration [45, 47, 48, 49], and the Bennett acceptance ratio method [50, 51]. The only limitation of our algorithm is that the Hamiltonian must be free of the sign problem, meaning that the weights in the partition function summation must be non-negative real numbers to allow a probabilistic interpretation—a common challenge for all QMC algorithms [52, 53].
@@
Rather than focusing on the full-state magic, we suggest that its volume-law corrections would be more important as they capture the nonlocal magic residing in correlations. Such kind of magic is spread nonlocally and cannot be removed via local non-Clifford operations. Notably, considering the volume-law corrections does not require a computable mixed state monotone and has similar effect to the mutual magic  [9, 54, 15, 55, 56], which we will detail in Sec. V.3. For both the 1D and 2D ground states of TFI models with finite system sizes, we observe discontinous indications of the volume-law corrections at the critical point, which reflects the sudden transition of the ground-state magical structure across the phase transition point.
@@ II.1 Short review of stabilizer protocols
$\mathcal{G}_{N}:=\bigg{\{}\xi\otimes_{j}\sigma_{j}\bigg{|}\sigma_{j}\in\{I,X,Y,Z\},\xi\in\{\pm 1,\pm i\}\bigg{\}}$ (2)
@@ II.2 $\alpha$-stabilizer Rényi entropy
$M_{\alpha}(\rho):=\frac{1}{1-\alpha}\log\bigg{[}\frac{1}{2^{N}}\sum_{P\in\mathcal{P}_{N}}{|\operatorname{Tr}(\rho P)|^{2\alpha}}\bigg{]}$ (3)
@@
The $2$-SRE $M_{2}(\rho)$ satisfies the following properties: (i) faithfulness: ${M}_{\alpha}(\rho)=0$ iff $\rho$ has no magic; stability: (ii) ${M}_{\alpha}(C\rho C^{\dagger})={M}_{\alpha}(\rho)$ for any Clifford unitary $C\in\mathcal{C}$ related to the Hilbert space; (iii) additivity: ${M}_{\alpha}(\rho\otimes\rho^{\prime})=M_{\alpha}(\rho)+M_{\alpha}(\rho^{\prime})$ for arbitary two pure states $\rho$ and $\rho^{\prime}$.
@@ II.3 Generalized partition functions
$Q:=\sum_{P\in\mathcal{P}_{N}}|\operatorname{Tr}(e^{-\beta H}P)|^{4}=\sum_{P\in\mathcal{P}_{N}}[\operatorname{Tr}(e^{-\beta H}P)]^{4}$ (7)
@@
$\begin{split}\tilde{M}_{2}(\rho)=&-\log\bigg{[}\frac{1}{2^{N}}\frac{Q}{Z^{4}}\bigg{]}+\log\bigg{[}\frac{Z_{2}}{Z^{2}}\bigg{]}\\ \end{split}$ (8)
@@ III.1 Stochastic series expansion
$\begin{split}\operatorname{Tr}(e^{-\beta H}P)=&\sum_{n}\frac{\beta^{n}}{n!}\operatorname{Tr}[(-H)^{n}P]\\ \end{split}$ (9)
@@
$\operatorname{Tr}(e^{-\beta H}P)=\sum_{n}\sum_{S_{n}}\frac{\beta^{n}}{n!}\operatorname{Tr}\bigg{\{}\bigg{[}\prod_{q=1}^{n}H_{\mu_{p},\nu_{p}}\bigg{]}P\bigg{\}}$ (10)
@@
$\operatorname{Tr}(e^{-\beta H}P)=\sum_{\alpha}\sum_{S_{\Lambda}}W(P;{S}_{\Lambda},\ket{\alpha_{0}})$ (11)
@@
Figure 1: (a) A configuration of $\operatorname{Tr}(e^{-\beta H}P)$ is shown in the language of SSE. The vertical axis represents the spatial degrees of freedom of the system, such as lattice sites $s_{0}$, $s_{1}$$\cdots$. In this exmple, a qubit lies at each site, with empty and solid circles denoting $\ket{\uparrow}$ and $\ket{\downarrow}$, respectively. The horizontal axis indicates the $(\Lambda+1)$ imaginary-time slices. Associated to each slice, the state of qubits are represented by |$\ket{\alpha_{q}}$. Then, |$\ket{\alpha_{0}}$ is sequentially propagated by operators $P=\prod_{j}\sigma^{j}$, $H_{\mu_{1},\nu_{1}}$, $H_{\mu_{2},\nu_{2}}$, $\cdots$, and $H_{\mu_{\Lambda},\nu_{\Lambda}}$, evolving through the intermidiate states |$\ket{\alpha_{1}}$, |$\ket{\alpha_{2}}$, $\cdots$, until it returns to |$\ket{\alpha_{\Lambda+1}}\equiv\ket{\alpha_{0}}$. For diagonal (off-diagonal) $\sigma_{j}$ and $H_{\mu_{q},\nu_{q}}$, we denote them with empty (solid) quadrangles in the diagram. (b) A sampled configuration of $Q$ with four replicas of $\operatorname{Tr}(e^{-\beta H}P)$. The four replicas share the same $P$, while |$\ket{\alpha_{0}^{k}}$ and $S_{\Lambda}^{k}$ can be different. Here we only show the time slice of $P$ for simplicity and the others are represented by $S_{\Lambda}^{k}$.
@@
$Q=\sum_{P\in\mathcal{P}_{N}}\sum_{\alpha^{k}}\sum_{{S}^{k}_{\Lambda}}\bigg{[}\prod_{k=1}^{4}W(P;{S}_{\Lambda}^{k},\ket{\alpha_{0}^{k}})\bigg{]}$ (13)
@@ III.2 Reduced configuration space
$Q=\sum_{\tilde{P}\in\tilde{\mathcal{P}}_{N}}\sum_{\tilde{\alpha}^{k}}\sum_{\tilde{S}^{k}_{\Lambda}}\bigg{[}\prod_{k=1}^{4}W(\tilde{P};\tilde{S}_{\Lambda}^{k},\ket{\tilde{\alpha}_{0}^{k}})\bigg{]}$ (16)
@@ IV.1 Reweight-annealing and importance sampling
$\begin{split}&\tilde{M}_{2}(\lambda)=\log\frac{Q(\lambda_{0})}{Q(\lambda)}-2\log\frac{Z(\lambda_{0})}{Z(\lambda)}-\log\frac{Z_{2}(\lambda_{0})}{Z_{2}(\lambda)}\end{split}$ (18)
Since the distribution overlap between $Q(\lambda_{0})$ and $Q(\lambda)$ (similar for $Z$ and $Z_{2}$) can be significantly small, practically, to enhance the simulation efficiency, interpolations are needed. We consdier
$\begin{split}&\frac{Q(\lambda_{0})}{Q(\lambda)}=\prod_{k=1}^{m}\frac{Q(\lambda_{k-1})}{Q(\lambda_{k})}\\ &\frac{Z(\lambda_{0})}{Z(\lambda)}=\prod_{k=1}^{m}\frac{Z(\lambda_{k-1})}{Z(\lambda_{k})}\\ &\frac{Z_{2}(\lambda_{0})}{Z_{2}(\lambda)}=\prod_{k=1}^{m}\frac{Z_{2}(\lambda_{k-1})}{Z_{2}(\lambda_{k})}\end{split}$ (19)
@@
$\begin{split}&\frac{Q(\beta_{k-1})}{Q(\beta_{k})}=\bigg{\langle}\bigg{(}\frac{\beta_{k-1}}{\beta_{k}}\bigg{)}^{n_{\text{tot}}}\bigg{\rangle}_{\beta_{k},Q}\\ &\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\bigg{\langle}\bigg{(}\frac{\beta_{k-1}}{\beta_{k}}\bigg{)}^{n_{\text{tot}}}\bigg{\rangle}_{\beta_{k},Z}\\ &\frac{Z_{2}(\beta_{k-1})}{Z_{2}(\beta_{k})}=\bigg{\langle}\bigg{(}\frac{\beta_{k-1}}{\beta_{k}}\bigg{)}^{n_{\text{tot}}}\bigg{\rangle}_{\beta_{k},Z_{2}}\\ \end{split}$ (20)
@@
$\begin{split}&\frac{Q(J_{k-1})}{Q(J_{k})}=\bigg{\langle}\bigg{(}\frac{J_{k-1}}{J_{k}}\bigg{)}^{n_{J,\text{tot}}}\bigg{\rangle}_{J_{k},Q}\\ &\frac{Z(J_{k-1})}{Z(J_{k})}=\bigg{\langle}\bigg{(}\frac{J_{k-1}}{J_{k}}\bigg{)}^{n_{J,\text{tot}}}\bigg{\rangle}_{J_{k},Z}\\ &\frac{Z_{2}(J_{k-1})}{Z_{2}(J_{k})}=\bigg{\langle}\bigg{(}\frac{J_{k-1}}{J_{k}}\bigg{)}^{n_{J,\text{tot}}}\bigg{\rangle}_{J_{k},Z_{2}}\\ \end{split}$ (21)

### v3

@@ Abstract
We present a novel quantum Monte Carlo method for evaluating the $\alpha$-stabilizer Rényi entropy (SRE) for any integer $\alpha\geq 2$. By interpreting $\alpha$-SRE as partition function ratios, we eliminate the sign problem in the imaginary-time path integral by sampling reduced Pauli strings within a reduced configuration space, which enables efficient classical computations of $\alpha$-SRE and its derivatives to explore magic in previously inaccessible 2D/higher-dimensional systems. We first isolate the free energy part in $2$-SRE, which is a trivial term. Notably, at quantum critical points in 1D/2D transverse field Ising (TFI) models, we reveal nontrivial singularities associated with the characteristic function contribution, directly tied to magic. Their interplay leads to complicated behaviors of $2$-SRE, avoiding extrema at critical points generally. In contrast, analyzing the volume-law correction to SRE reveals a discontinuity tied to criticalities, suggesting that it is more informative than the full-state magic. For conformal critical points, we claim it could reflect nonlocal magic residing in correlations. Finally, we verify that $2$-SRE fails to characterize magic in mixed states (e.g. Gibbs states), yielding nonphysical results. This work provides a powerful tool for exploring the roles of magic in large-scale many-body systems, and reveals intrinsic relation between magic and many-body physics.
@@ I Introduction
Characterizing many-body behaviors in quantum systems has been a central task since the last century [1, 2]. Despite the locality of interactions in most quantum matters, the intricate correlations that give rise to rich and exotic collective phenomena present formidable theoretical and computational hurdles for physicists. During the past decades, remarkable headways are made by introducing the theory of quantum information into many-body physics. Quantum entanglement, with its quantum nature and as probably the most representative quantum resource [3], has turned out to relate closely to phase transitions, conformal field theories and topological orders [4, 5, 6].
In recent years, another quantum resource called magic or non-stabilizerness, which is closely associated with the stabilizer formalism [7, 8], has also entered the field of vision in many-body physics [9, 10, 11]. According to the celebrated Gottesman-Knill theorem [12, 13, 14], magic, rather than merely entanglement, is what necessarily enables quantum computation to outperform classical computation, as simulating a highly entangled stabilizer state requires only polynomial resources on a classical computer with Clifford protocols.
@@
The contributions of this work span both algorithmic and physical levels. At the algorithmic level, we introduce the first unbiased Quantum Monte Carlo (QMC) algorithm capable of computing both the $\alpha$-SRE and its derivatives in large-scale and high-dimensional quantum many-body systems within polynomial time, as long as the Hamiltonian is sign-problem free. In fact, over the past years, several classical numerical tools have been developed to calculate the $\alpha$-SRE including the methods with matrix product states (MPS) and tree tensor networks [16, 17, 18, 11, 19, 35]. However, these methods face limitations in two or higher dimensions and finite-temperature cases. A recent hybrid algorithm proposed in Ref. [36] addresses this issue by expressing the $\alpha$-SRE using the language of tensor network (TN) and then applying non-equilibrium QMC sampling [37]. However, their algorithm computes $\alpha$-SRE at only a single point per run and is incapable of extracting derivative information, yielding results with limited physical insight.
Technically, we prove that the QMC simulations of the $\alpha$-SRE, which relates to some generalized partition function, can be restricted in the reduced configuration space by sampling reduced Pauli strings (both terms are defined in Sec. III). Using this algorithm, it is straightforward to estimate the values and derivatives of $\alpha$-SRE by considering the partition function difference, in conjunction with Monte Carlo techniques including but not restricted to: the reweight-annealing method (importance sampling) [38, 39, 40, 41, 42, 43, 44, 45, 46], thermodynamic integration [45, 47, 48, 49], and the Bennett acceptance ratio method [50, 51]. The flexibility of our framework also allows for the integration of ideas from recent QMC algorithms for computing entanglement entropy, which could lead to further performance improvements [37, 52, 53, 54, 55, 56, 57, 58]. The only limitation of our algorithm is that the Hamiltonian must be free of the sign problem, meaning that the weights in the partition function summation must be non-negative real numbers to allow a probabilistic interpretation—a common challenge for all QMC algorithms [59, 60]. However, sign-problem-free models already host a rich variety of fundamental many-body phenomena, including spontaneous symmetry breaking, topological phase transitions, topological order, and conformal criticality. As such, our method plays a crucial role in advancing the understanding of quantum magic and complex quantum matter.
@@
Rather than focusing on the full-state magic, we suggest that its volume-law corrections would be more important as a nonzero volume-law correction is a necessary condition for the existence of nonlocal magic residing in correlations. Such kind of magic is spread nonlocally and cannot be removed via local non-Clifford operations. Notably, considering the volume-law corrections does not require a computable mixed state monotone and has similar effect to the mutual magic  [9, 61, 15, 62, 63], which we will detail in Sec. V.3. For both the 1D and 2D ground states of TFI models with finite system sizes, we observe discontinuous indications of the volume-law corrections at the critical point, which reflects the sudden transition of the ground-state magical structure across the phase transition point.
@@ II.1 Short review of stabilizer protocols
$\mathcal{G}_{N}:=\bigg{\{}\xi\otimes_{j}\sigma_{j}\bigg{|}\sigma_{j}\in\{I,X,Y,Z\},\xi\in\{\pm 1,\pm i\}\bigg{\}},$ (2)
@@ II.2 $\alpha$-stabilizer Rényi entropy
$M_{\alpha}(\rho):=\frac{1}{1-\alpha}\log\bigg{[}\frac{1}{2^{N}}\sum_{P\in\mathcal{P}_{N}}{|\operatorname{Tr}(\rho P)|^{2\alpha}}\bigg{]},$ (3)
@@
The $2$-SRE $M_{2}(\rho)$ satisfies the following properties: (i) faithfulness: ${M}_{\alpha}(\rho)=0$ iff $\rho$ has no magic; (ii) stability: ${M}_{\alpha}(C\rho C^{\dagger})={M}_{\alpha}(\rho)$ for any Clifford unitary $C\in\mathcal{C}$ related to the Hilbert space; (iii) additivity: ${M}_{\alpha}(\rho\otimes\rho^{\prime})=M_{\alpha}(\rho)+M_{\alpha}(\rho^{\prime})$ for arbitrary two pure states $\rho$ and $\rho^{\prime}$.
@@ II.3 Generalized partition functions
$Q:=\sum_{P\in\mathcal{P}_{N}}|\operatorname{Tr}(e^{-\beta H}P)|^{4}=\sum_{P\in\mathcal{P}_{N}}[\operatorname{Tr}(e^{-\beta H}P)]^{4},$ (7)
@@
$\begin{split}\tilde{M}_{2}(\rho)=&-\log\bigg{[}\frac{1}{2^{N}}\frac{Q}{Z^{4}}\bigg{]}+\log\bigg{[}\frac{Z_{2}}{Z^{2}}\bigg{]}.\end{split}$ (8)
@@ III.1 Stochastic series expansion
$\begin{split}\operatorname{Tr}(e^{-\beta H}P)=&\sum_{n}\frac{\beta^{n}}{n!}\operatorname{Tr}[(-H)^{n}P].\end{split}$ (9)
@@
$\operatorname{Tr}(e^{-\beta H}P)=\sum_{n}\sum_{S_{n}}\frac{\beta^{n}}{n!}\operatorname{Tr}\bigg{\{}\bigg{[}\prod_{q=1}^{n}H_{\mu_{p},\nu_{p}}\bigg{]}P\bigg{\}},$ (10)
@@
$\operatorname{Tr}(e^{-\beta H}P)=\sum_{\alpha}\sum_{S_{\Lambda}}W(P;{S}_{\Lambda},\ket{\alpha_{0}}),$ (11)
@@
Figure 1: (a) A configuration of $\operatorname{Tr}(e^{-\beta H}P)$ is shown in the language of SSE. The vertical axis represents the spatial degrees of freedom of the system, such as lattice sites $s_{0}$, $s_{1}$$\cdots$. In this example, a qubit lies at each site, with empty and solid circles denoting $\ket{\uparrow}$ and $\ket{\downarrow}$, respectively. The horizontal axis indicates the $(\Lambda+1)$ imaginary-time slices. Associated to each slice, the state of qubits are represented by |$\ket{\alpha_{q}}$. Then, |$\ket{\alpha_{0}}$ is sequentially propagated by operators $P=\prod_{j}\sigma^{j}$, $H_{\mu_{1},\nu_{1}}$, $H_{\mu_{2},\nu_{2}}$, $\cdots$, and $H_{\mu_{\Lambda},\nu_{\Lambda}}$, evolving through the intermediate states |$\ket{\alpha_{1}}$, |$\ket{\alpha_{2}}$, $\cdots$, until it returns to |$\ket{\alpha_{\Lambda+1}}\equiv\ket{\alpha_{0}}$. For diagonal (off-diagonal) $\sigma_{j}$ and $H_{\mu_{q},\nu_{q}}$, we denote them with empty (solid) quadrangles in the diagram. (b) A sampled configuration of $Q$ with four replicas of $\operatorname{Tr}(e^{-\beta H}P)$. The four replicas share the same $P$, while |$\ket{\alpha_{0}^{k}}$ and $S_{\Lambda}^{k}$ can be different. Here we only show the time slice of $P$ for simplicity and the others are represented by $S_{\Lambda}^{k}$.
@@
$Q=\sum_{P\in\mathcal{P}_{N}}\sum_{\alpha^{k}}\sum_{{S}^{k}_{\Lambda}}\bigg{[}\prod_{k=1}^{4}W(P;{S}_{\Lambda}^{k},\ket{\alpha_{0}^{k}})\bigg{]},$ (13)
@@ III.2 Reduced configuration space
$Q=\sum_{\tilde{P}\in\tilde{\mathcal{P}}_{N}}\sum_{\tilde{\alpha}^{k}}\sum_{\tilde{S}^{k}_{\Lambda}}\bigg{[}\prod_{k=1}^{4}W(\tilde{P};\tilde{S}_{\Lambda}^{k},\ket{\tilde{\alpha}_{0}^{k}})\bigg{]},$ (16)
@@ IV.1 Reweight-annealing and importance sampling
$\begin{split}&\tilde{M}_{2}(\lambda)=\log\frac{Q(\lambda_{0})}{Q(\lambda)}-2\log\frac{Z(\lambda_{0})}{Z(\lambda)}-\log\frac{Z_{2}(\lambda_{0})}{Z_{2}(\lambda)}.\end{split}$ (18)
Since the distribution overlap between $Q(\lambda_{0})$ and $Q(\lambda)$ (similar for $Z$ and $Z_{2}$) can be significantly small, practically, to enhance the simulation efficiency, interpolations are needed. We consider
$\begin{split}&\frac{Q(\lambda_{0})}{Q(\lambda)}=\prod_{k=1}^{m}\frac{Q(\lambda_{k-1})}{Q(\lambda_{k})},\\ &\frac{Z(\lambda_{0})}{Z(\lambda)}=\prod_{k=1}^{m}\frac{Z(\lambda_{k-1})}{Z(\lambda_{k})},\\ &\frac{Z_{2}(\lambda_{0})}{Z_{2}(\lambda)}=\prod_{k=1}^{m}\frac{Z_{2}(\lambda_{k-1})}{Z_{2}(\lambda_{k})},\end{split}$ (19)
@@
$\begin{split}&\frac{Q(\beta_{k-1})}{Q(\beta_{k})}=\bigg{\langle}\bigg{(}\frac{\beta_{k-1}}{\beta_{k}}\bigg{)}^{n_{\text{tot}}}\bigg{\rangle}_{\beta_{k},Q},\\ &\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\bigg{\langle}\bigg{(}\frac{\beta_{k-1}}{\beta_{k}}\bigg{)}^{n_{\text{tot}}}\bigg{\rangle}_{\beta_{k},Z},\\ &\frac{Z_{2}(\beta_{k-1})}{Z_{2}(\beta_{k})}=\bigg{\langle}\bigg{(}\frac{\beta_{k-1}}{\beta_{k}}\bigg{)}^{n_{\text{tot}}}\bigg{\rangle}_{\beta_{k},Z_{2}},\end{split}$ (20)
@@
$\begin{split}&\frac{Q(J_{k-1})}{Q(J_{k})}=\bigg{\langle}\bigg{(}\frac{J_{k-1}}{J_{k}}\bigg{)}^{n_{J,\text{tot}}}\bigg{\rangle}_{J_{k},Q},\\ &\frac{Z(J_{k-1})}{Z(J_{k})}=\bigg{\langle}\bigg{(}\frac{J_{k-1}}{J_{k}}\bigg{)}^{n_{J,\text{tot}}}\bigg{\rangle}_{J_{k},Z},\\ &\frac{Z_{2}(J_{k-1})}{Z_{2}(J_{k})}=\bigg{\langle}\bigg{(}\frac{J_{k-1}}{J_{k}}\bigg{)}^{n_{J,\text{tot}}}\bigg{\rangle}_{J_{k},Z_{2}},\end{split}$ (21)